# Sports-QA: A Large-Scale Video Question Answering Benchmark for Complex and Professional Sports

### v2

@@ Abstract
Reasoning over sports videos for question answering is an important task with numerous applications, such as player training and information retrieval. However, this task has not been explored due to the lack of relevant datasets and the challenging nature it presents. Most datasets for video question answering (VideoQA) focus mainly on general and coarse-grained understanding of daily-life videos, which is not applicable to sports scenarios requiring professional action understanding and fine-grained motion analysis. In this paper, we introduce the first dataset, named Sports-QA, specifically designed for the sports VideoQA task. The Sports-QA dataset includes various types of questions, such as descriptions, chronologies, causalities, and counterfactual conditions, covering multiple sports. Furthermore, to address the characteristics of the sports VideoQA task, we propose a new Auto-Focus Transformer (AFT) capable of automatically focusing on particular scales of temporal information for question answering. We conduct extensive experiments on Sports-QA, including baseline studies and the evaluation of different methods. The results demonstrate that our AFT achieves state-of-the-art performance [^1] Project homepage: https://github.com/HopLee6/Sports-QA.
@@ 1 Introduction
Sports video analysis has been attracting increasing attention in recent years [39, 18, 17, 42, 22, 33]. While research progress has been made on tasks such as sports action recognition [19, 42, 25], reasoning over sports videos for question answering has not been explored. As humans, we can not only recognize the actions of the players in sports videos but also understand the effects of the players’ actions, explain why a team loses the score, and imagine what would happen under counterfactual situations. Our impressive capabilities in reasoning allow us to answer complex questions related to sports videos, which is crucial in applications such as obtaining crucial statistics in matches for player/team performance qualification, analyzing players’ actions and team strategies for performance improvement, and efficiently retrieving information for audiences and analysts. While sports video reasoning is clearly important, it is under-explored due to challenges and the lack of datasets.
@@
In this paper, we tackle sports VideoQA, focusing on reasoning over sports videos for question answering. This is distinct from general VideoQA, which centers on general video understanding, mainly encompassing descriptive and temporal aspects [14, 36, 34, 38], as depicted in Figure 3. Compared to general VideoQA, sports VideoQA is more challenging because the questions may involve particular teams and professional actions, and it requires a fine-grained understanding of actions and intra/inter-team relation modeling. As pointed out in [28]: 1) coarse-grained global video understanding is insufficient for sports analysis where the actions are rapid and hard to differentiate; 2) the models pre-trained on coarse-grained datasets are not always beneficial to fine-grained tasks. A large-scale dataset containing professional sports scenarios with detailed annotations of sports-relevant questions is thus required for explicit and comprehensive reasoning for sports VideoQA.
@@
Table 1: VideoQA datasets comparison. “Auto/Man” is “automatic/manual”. “MC/OE” is “multi-choice/open-ended”.
@@
  • We contribute a large-scale dataset, Sports-QA, which consists of various types of questions and multiple sports for VideoQA. To our knowledge, it is the first dataset for complex reasoning over professional sports actions.
@@
  • We conduct extensive experiments on our dataset, including baseline study, evaluation of existing methods, and visualization of predictions. The results show the superiority of the proposed AFT for the sports VideoQA task.
@@ 2 Related Work
VideoQA Datasets. The development of VideoQA has been greatly facilitated by the emergence of various datasets, such as TGIF-QA [14], MSVD-QA, and MSRVTT-QA [36], DramaQA [4], and NExT-QA [34]. TGIF-QA, for instance, offers a comprehensive set of four sub-tasks designed for temporal reasoning in general videos and is widely recognized in the VideoQA community. MSVD-QA and MSRVTT-QA are open-ended datasets constructed from existing video captions, while DramaQA focuses on understanding drama stories with hierarchical QAs and character-centered video annotation. NExT-QA, introduced for describing and explaining temporal actions, provides both multi-choice and open-ended questions through manual annotation. However, existing datasets predominantly center around general video understanding in daily scenarios or involve coarse-grained action/event reasoning. In contrast, our contribution lies in fine-grained and professional analysis within sports scenarios. We present Sports-QA, a dataset that addresses aspects of description, chronology, causality, and counterfactual conditions for multiple sports.
VideoQA Methods. VideoQA poses a significant challenge as it necessitates models to grasp both spatial and temporal information from videos to answer questions. Various deep models have been developed, approaching this task from different perspectives [7, 15, 20, 1, 9]. For instance, the deep heterogeneous graph alignment network by Jiang et al. [15] addresses VideoQA by simultaneously aligning intra/inter-modality information. Another approach involves a multimodal attention model proposed by Fan et al. [7], where heterogeneous memory learns global context from visual features, and question memory captures the complex semantics of questions. IGV, introduced by Li et al. [20], grounds question-critical scenes in videos by considering causal relations that remain invariant to complement contents. However, a limitation in these approaches is the lack of consideration for the fact that different questions may require temporal dependencies of specific scales. In response to this challenge, we propose the Auto-Focus Transformer, designed to automatically focus on a specific temporal scale based on the question for VideoQA.
@@ 3.1 Data Source
Regarding the collection of sports videos, we consider the following aspects: 1) The visual quality of the video data should be high enough to conduct fine-grained video understanding, such as video resolution and frame rate. 2) Instead of applying to a single type of sports, we expect a VideoQA dataset involving multiple sports. After a deep survey of the works on sports video understanding, we find that the MultiSports [19] and FineGym [28] datasets are highly suitable for our purpose. Specifically, MultiSports is a dataset for the detection of various sports actions, including four sports (i.e., basketball, football, volleyball, and aerobic gymnastics) and 66 fine-grained action categories. FineGym is a dataset for the recognition of fine-grained actions in gymnastics, covering four types (i.e., vault, floor exercise, uneven bars, and balance beam) and 530 sub-action categories. The action categories in MultiSports and FineGym are professional terms (such as “volleyball spike”, “football tackle”, “basketball defensive rebound”, and “vault Salto backward tucked”) instead of common and atomic actions (such as “run” and “stand”). Such professional labels are of great value to our dataset. For more details on MultiSports and FineGym, please refer to the supplementary material. Since the length of videos varies greatly (from a few seconds to a few minutes), we first segment the videos into clips and then generate QA pairs based on these clips.
@@ 3.2.1 QA Pair Generation for Ball Games
Specifically, for ball games, including basketball, football, and volleyball, we define five attributes (Team, Outcome, Cause of Outcome, Cause of Action, and Effect of Action) for actions. We then manually annotate the attributes of each action, and the definitions and annotation process of these attributes are elaborated in the supplementary material. Based on the attributes, we generate questions about the videos by designing various templates (listed in the supplementary material) to inquire about these attributes. For instance, given a video shown in Figure 2, after annotating the attributes for the action “spike”, a question querying a particular attribute (e.g., Outcome) of the action can be generated, such as “Is the ‘spike’ of the right team successful?”. Specifically, our dataset involves four types of questions: descriptive, temporal, causal, and counterfactual. The details of each type are described as follows.
@@
1. Descriptive: Descriptive questions require holistic comprehension and involve various aspects of information from the videos. This type of question includes simple queries such as “What is the video about?” and “Does SOME-TEAM perform SOME-ACTION?”, as well as complex ones involving counting and the outcome of actions, such as “How many times does SOME-TEAM perform SOME-ACTION?” and “Does SOME-TEAM successfully do their i-th SOME-ACTION?
@@
In particular, the questions are queries about the attributes, and their answers can be obtained directly from our annotations or through logical reasoning over the annotations. These answers include responses like “yes/no”, numerical values, sports names (such as “volleyball”), action categories in MultiSports and FineGym (such as “block”), and short phrases describing the reasons (such as “out of bounds”). See Figure 2 for more concrete examples of the attributes and the corresponding generated QA pairs. Additional examples are provided in the supplementary material.
@@
3.2.2 QA Pair Generation for Gymnastic Sport
@@
3.3 Data Processing
Unlike daily scenarios, the players are highly restricted by rules in sports. For example, after the left team does “volleyball serve”, the right team must do “first-pass”. Therefore, there exist highly correlated actions in the videos, and thus some questions and answers are highly correlated in our generated pairs. Such questions are meaningless to VideoQA because the answers can be easily inferred by rules or common sense, which should be removed. To this end, we first obtain the meta-questions by removing the team information and the order of actions from the questions. For example, the meta-question of “What does the left team do after the other team does the second spike?” is “What does the team do after the other team does spike?”. We then examine the correlations between the meta-questions and their answers. If there is only one answer to the meta-question, the QA pairs are removed; If an answer to a meta-question occurs more frequently than other answers (frequency larger than 0.5), the corresponding QA pairs are randomly removed to make the frequency lower than 0.5. Following [14, 36], we define an open-ended task based on the generated QA pairs. Specifically, all answers in the QA pairs form an answer pool, which is regarded as 191 classes (after discarding the answer classes that contain fewer than 30 samples).
@@ 3.3 Data Processing
Unlike daily scenarios, players in sports are highly restricted by rules. For instance, after the left team executes a “volleyball serve”, the right team must follow with a “first-pass”. This results in highly correlated actions in the videos, leading to questions and answers that are also highly correlated in our generated pairs. However, such questions are often meaningless for VideoQA, as the answers can be easily inferred through rules or common sense and should be removed. To address this, we first obtain meta-questions by removing team information and the order of actions from the original questions. For example, the meta-question for “What does the left team do after the other team does the second spike?” becomes “What does the team do after the other team does spike?” We then examine the correlations between the meta-questions and their answers. If there is only one answer to the meta-question, we remove the corresponding QA pairs. If an answer to a meta-question occurs more frequently than other answers (frequency larger than 0.5), we randomly remove the corresponding QA pairs to balance the frequencies, ensuring they are all lower than 0.5. Following the approach in [14, 36], we define an open-ended task based on the generated QA pairs. Specifically, all answers in the QA pairs form an answer pool, which is treated as 191 classes (after discarding answer classes with fewer than 30 samples).

### v4

@@ Abstract
Reasoning over sports videos for question answering is an important task with numerous applications, such as player training and information retrieval. However, this task has not been explored due to the lack of relevant datasets and the challenging nature it presents. Most datasets for video question answering (VideoQA) focus mainly on general and coarse-grained understanding of daily-life videos, which is not applicable to sports scenarios requiring professional action understanding and fine-grained motion analysis. In this paper, we introduce the first dataset, named Sports-QA, specifically designed for the sports VideoQA task. The Sports-QA dataset includes various types of questions, such as descriptions, chronologies, causalities, and counterfactual conditions, covering multiple sports. Furthermore, to address the characteristics of the sports VideoQA task, we propose a new Auto-Focus Transformer (AFT) capable of automatically focusing on particular scales of temporal information for question answering. We conduct extensive experiments on Sports-QA, including baseline studies and the evaluation of different methods. The results demonstrate that our AFT achieves state-of-the-art performance.
keywords:
Video Question Answering, Sports Video, Benchmark, Auto-Focus Transformer
@@ 1 Introduction
Sports video analysis has been attracting increasing attention in recent years [45, 19, 17, 49, 26, 38, 22]. While research progress has been made on tasks such as sports action recognition [18, 49, 29], reasoning over sports videos for question answering has not been explored. As humans, we can not only recognize the actions of the players in sports videos but also understand the effects of the players’ actions, explain why a team loses the score, and imagine what would happen under counterfactual situations. Our impressive capabilities in reasoning allow us to answer complex questions related to sports videos, which is crucial in applications such as obtaining crucial statistics in matches for player/team performance qualification, analyzing players’ actions and team strategies for performance improvement, and efficiently retrieving information for audiences and analysts. While sports video reasoning is clearly important, it is under-explored due to challenges and the lack of datasets.
@@
In this paper, we tackle sports VideoQA, focusing on reasoning over sports videos for question answering. This is distinct from general VideoQA, which centers on general video understanding, mainly encompassing descriptive and temporal aspects [15, 41, 39, 46], as depicted in Fig. 2. Compared to general VideoQA, sports VideoQA is more challenging because the questions may involve particular teams and professional actions, and it requires a fine-grained understanding of actions and intra/inter-team relation modeling. As pointed out in [33]: 1) coarse-grained global video understanding is insufficient for sports analysis where the actions are rapid and hard to differentiate; 2) the models pre-trained on coarse-grained datasets are not always beneficial to fine-grained tasks. In the meantime, although significant strides have been made in the fine-grained recognition of individual actions [33, 23, 32], these datasets may fall short in effectively evaluating the ability for temporal and causal reasoning over actions performed by individuals or teams in sports scenarios. A large-scale dataset containing professional sports scenarios with detailed annotations of sports-relevant questions is thus required for explicit and comprehensive reasoning for sports VideoQA.
@@
Table 1: VideoQA datasets comparison. “Auto/Man” represents “automatic/manual”. “MC/OE” represents “multi-choice/open-ended”.
@@
  • We contribute a large-scale dataset, which consists of various types of questions and multiple sports for VideoQA. To our knowledge, it is the first dataset for complex reasoning over professional sports actions.
@@
  • We conduct extensive experiments on our dataset, including baseline study, evaluation of existing methods, and visualization of predictions. The results show the superiority of the proposed AFT for sports VideoQA.
@@ 2 Related Work
VideoQA Datasets. The development of VideoQA has been greatly facilitated by the emergence of various datasets, such as TGIF-QA [15], MSVD-QA, and MSRVTT-QA [41], DramaQA [3], NExT-QA [39], NExT-OOD [50]. TGIF-QA, for instance, offers a comprehensive set of four sub-tasks designed for temporal reasoning in general videos and is widely recognized in the VideoQA community. MSVD-QA and MSRVTT-QA are open-ended datasets constructed from existing video captions, while DramaQA focuses on understanding drama stories with hierarchical QAs and character-centered video annotation. NExT-QA, introduced for describing and explaining temporal actions, provides both multi-choice and open-ended questions through manual annotation. However, existing datasets predominantly center around general video understanding in daily scenarios or involve coarse-grained action/event reasoning. In contrast, our contribution lies in fine-grained and professional analysis within sports scenarios. We present Sports-QA, a dataset that addresses aspects of description, chronology, causality, and counterfactual conditions for multiple sports of diverse characteristics.
VideoQA Methods. VideoQA poses a significant challenge as it necessitates models to grasp both spatial and temporal information from videos to answer questions. Various deep models have been developed, approaching this task from different perspectives [8, 14, 24, 1, 10, 42, 44, 25, 20]. For instance, the deep heterogeneous graph alignment network by Jiang et al. [14] addresses VideoQA by simultaneously aligning intra/inter-modality information. Another approach involves a multimodal attention model proposed by Fan et al. [8], where heterogeneous memory learns global context from visual features, and question memory captures the complex semantics of questions. IGV, introduced by Li et al. [24], grounds question-critical scenes in videos by considering causal relations that remain invariant to complement contents. However, a limitation in these approaches is the lack of consideration for the fact that different questions may require temporal dependencies of specific scales. In response to this challenge, we propose the Auto-Focus Transformer, designed to automatically focus on a specific temporal scale based on the question for VideoQA.
@@ 3 Sports-QA Dataset
Table 2: The statistics for the MultiSports dataset are presented.
Sports # Action # Instance Avg. Action/Video Duration # Bounding box
Gym 21 8,703 1.5s / 30.7s 325K
Volleyball 12 7,645 0.7s / 10.5s 139K
Football 15 12,254 0.7s / 22.6s 225K
Basketball 18 9,099 0.9s / 19.7s 213K
Total 66 37,701 1.0s / 20.9s 902K
[…]
@@ 3.1 Data Source
Regarding the collection of sports videos, we consider the following aspects: 1) The visual quality of the video data should be high enough to conduct fine-grained video understanding, such as video resolution and frame rate. 2) Instead of applying to a single type of sports, we expect a VideoQA dataset involving multiple sports. After a deep survey of the works on sports video understanding, we find that the MultiSports [18] and FineGym [33] datasets are highly suitable for our purpose. The details of MultiSports and FineGym are as follows.
MultiSports [18] is a dataset for the temporal localization of sports actions, encompassing four sports (i.e., basketball, football, volleyball, and aerobic gymnastics) and 66 fine-grained action categories [^3] Project homepage: https://deeperaction.github.io/datasets/multisports.html. The action categories in MultiSports are professional terms (such as “volleyball spike”, “football tackle”, and “basketball defensive rebound”) instead of common and atomic actions like “run” and “stand”. For each action instance, the dataset provides the bounding boxes of a player from the starting frame of an action to the ending frame, forming the action spatial-temporal tube. The statistics of MultiSports are shown in Table 2.
FineGym [33] is a dataset designed to elevate the field of action recognition by addressing the limitations observed in existing techniques [^4] Project homepage: https://sdolivia.github.io/FineGym. Developed to surpass current benchmarks, it offers a unique combination of richness, quality, and diversity in its content. This dataset is constructed on gymnasium videos, providing a realistic and varied environment for action recognition studies. What sets FineGym apart from other datasets is its meticulous temporal annotation at both action and sub-action levels, featuring a three-level semantic hierarchy. This hierarchical structure allows for a more nuanced understanding of activities, enabling researchers to explore and analyze actions in finer detail. As an illustrative example, an event is annotated as a sequence of elementary sub-actions Importantly, each sub-action within these sets is further annotated with finely defined class labels, contributing to a higher level of granularity in action recognition. Fig. 2 shows the action hierarchy of the MultiSports and FineGym datasets.
Figure 2: The action hierarchy of the MultiSports dataset is depicted at the top, while the FineGym dataset’s hierarchy is shown at the bottom. It’s important to note that the figure includes only four example actions for each sport.
The action categories in MultiSports and FineGym are professional terms (such as “volleyball spike”, “football tackle”, “basketball defensive rebound”, and “vault Salto backward tucked”) instead of common and atomic actions (such as “run” and “stand”). Such professional labels are of great value to our dataset. Since the length of videos varies greatly (from a few seconds to a few minutes), we segment the videos into clips and generate QA pairs based on these clips.
@@ 3.2.1 QA Pair Generation for Ball Games
Specifically, for ball games, including basketball, football, and volleyball, we define five attributes (Team, Outcome, Cause of Outcome, Cause of Action, and Effect of Action) for actions. We then manually annotate the attributes of each action. The definitions and annotation process of these attributes are elaborated as follows.
Attribute Definition. MultiSports consists of 45 action categories from ball games, with some being crucial for professional sports statistics or quantifying the performance of players/teams. For instance, the “2-Point Field Goal Percentage” in a basketball match necessitates a fine-grained understanding of all “2-point shot” actions, while the “saving” actions in football directly reflect the performance of goalkeepers. In this work, we focus on 28 crucial actions, listed in Table LABEL:acts. Considering both practical applications and research purposes, we define five attributes for each of the crucial action instances as follows.
1. Team: The team that the player of the action belongs to is denoted by two options: left team or right team. Specifically, in basketball and football, the determination of the left or right team depends on the team attacking towards the right or left. In volleyball, the left or right team is designated based on their position relative to the left or right side of the net.
2. Outcome: This attribute assesses whether the result of the action meets the expectation and is defined as a binary label: successful or failed. For example, if a “2-point shot” in basketball scores a goal, it is annotated as successful; otherwise, it is labeled as failed.
3. Cause of Outcome: This attribute indicates the cause of failure for offensive actions, such as a “2-point shot” in basketball, “shoot” in football, and “spike” in volleyball, or the cause of success for defensive actions, like “block” in football and “save” in volleyball. Specifically, the cause of failure for offensive actions reflects player shortcomings. For example, the failure of an offensive “long pass” in football could be attributed to defensive interception, bad pass, or bad catch. Identifying the precise cause helps teams identify their weaknesses and improve performance. Similarly, the cause of success for defensive actions highlights player strengths. For instance, the success of a defensive “save” in volleyball might be due to offensive out of bounds, offensive blocked by net, or actually catch the ball, with the last case directly reflecting defensive performance. We do not consider the cause of success for offensive actions or the cause of failure for defensive actions, as they cannot be attributed to different cases. For example, the failure of “football saving” can only be explained by the goalkeeper missing the ball. Therefore, these causes are not worth discussing. We define various cause options for applicable action categories and ask annotators to choose from them. Note that this attribute is applicable to certain action categories, and the causes vary for different action categories.
4. Cause of Action: This attribute indicates the actions that cause the current action. Each action instance in a video is initially labeled with its time order, based on the start time, which serves as the unique ID of the action in the video. For each crucial action, annotators are asked to provide the IDs of the actions that cause the current action; these causes are not necessarily crucial actions. Consequently, the Cause of Action is a list consisting of the IDs of actions that are mostly temporally adjacent to the current action. Not all crucial action categories are required to provide this attribute because the causes of some actions are unique. For instance, the only cause of “volleyball first pass” is “volleyball serve”.
5. Effect of Action: This attribute aims to identify the actions caused by the current action, analogous to Cause of Action. Although the cause and effect are conjunctive, we may consider only one of them, as some actions are not crucial. For instance, if ACTION M (crucial action) causes ACTION N (non-crucial action), we only label ACTION N as an effect of ACTION M.
Attribute Annotation. After defining the attributes mentioned above, we proceed to annotate them for each crucial action instance in MultiSports. The annotation process is divided into three stages: pre-annotation, formal annotation, and quality check. In the pre-annotation stage, annotators are grouped into three categories, with each group assigned responsibility for a specific type of ball game. Each annotator within a group is then assigned several videos corresponding to their designated sport. During this stage, annotators are tasked with labeling the attributes of the first 50 action instances. Following this, an intra-group check is conducted to address any issues related to understanding bias or potential mistakes, as well as to handle unexpected or rare situations. The annotation can only progress to the next stage when annotators achieve a consistent understanding of each attribute, as outlined in the protocols. Moving to the formal annotation stage, annotators use shared protocols to amend their previous annotations and label the remaining action instances. Once all actions are labeled, we perform an inter-group quality check on all annotations. The purpose of this inter-group check is to ensure that the annotations align with the common understanding of average individuals. The annotation process is completed within one month and involves the collaboration of 15 graduate students.
Based on the attributes, we generate questions about the videos by designing various templates (listed in the end of the paper) to inquire about these attributes. For instance, given a video shown in Fig. 3, after annotating the attributes for the action “spike”, a question querying a particular attribute (e.g., Outcome) of the action can be generated, such as “Is the ‘spike’ of the right team successful?”. Specifically, our dataset involves four types of questions: descriptive, temporal, causal, and counterfactual. The details of each type are described as follows.
@@
1. Descriptive: Descriptive questions require holistic comprehension and involve various aspects of information from the videos. These questions include simple queries such as “What is the video about?” and “Does SOME-TEAM perform SOME-ACTION?”, as well as complex ones involving counting and the outcome of actions, such as “How many times does SOME-TEAM perform SOME-ACTION?” and “Does SOME-TEAM successfully do their i-th SOME-ACTION?
@@
From Attributes to Question-Answer Pairs. In particular, the questions are queries about the attributes, and their answers can be obtained directly from our annotations or through logical reasoning and statistical analysis over the annotations. For instance, for a question asking about the cause of a specific action, the answer can be easily retrieved from its attribute of cause; For a question asking whether a specific action is successful, we can check its outcome attribute. We guarantee the correctness of the answers as we have checked the annotations of attributes as mentioned previously. These answers include responses like “yes/no”, numerical values, sports names (such as “volleyball”), action categories in MultiSports and FineGym (such as “block”), and short phrases describing the reasons (such as “out of bounds”). See Fig. 3 for more concrete examples of the attributes and the corresponding generated QA pairs.
@@
3.2.2 QA Pair Generation for Gymnastics
@@
3.3 Diversity, Debiasing and Problem Setting
Diversity of Questions. The limited diversity in generating questions using pre-designed templates may not significantly impact the significance of datasets for several reasons. First, in the context of sports, there is a high level of restriction imposed by rules and specific content, resulting in a more focused attention from the audience on particular actions and events. This inherent structure in sports content narrows down the range of potential questions, making template-based question generation a feasible approach. Secondly, the practice of template-based question generation is widely accepted and employed in the construction of VideoQA datasets [15, 41, 43]. The adoption of this approach in reputable datasets attests to its effectiveness and acceptance. Therefore, while the questions are template-based, their relevance and significance in the context of sports are preserved, ensuring the robustness of the dataset for analysis and evaluation.
@@ 3.3 Diversity, Debiasing and Problem Setting
Debiasing. Unlike daily scenarios, players in sports are highly restricted by rules. For instance, after the left team executes a “volleyball serve”, the right team must follow with a “first-pass”. This results in highly correlated actions in the videos, leading to questions and answers that are also highly correlated in our generated pairs. However, such questions are often meaningless for VideoQA, as the answers can be easily inferred through rules or common sense and should be removed. To address this, we first obtain meta-questions by removing team information and the order of actions from the original questions. For example, the meta-question for “What does the left team do after the other team does the second spike?” becomes “What does the team do after the other team does spike?” We then examine the correlations between the meta-questions and their answers. If there is only one answer to the meta-question, we remove the corresponding QA pairs. If an answer to a meta-question occurs more frequently than other answers (frequency larger than 0.5), we randomly remove the corresponding QA pairs to balance the frequencies, ensuring they are all lower than 0.5.
Problem Setting. Following the approach in [15, 41], we define an open-ended task based on the generated QA pairs. It is important to note that the open-ended setting, along with the multiple-choice setting, is widely used in the field of VideoQA. We have chosen to commit to the open-ended task instead of the multiple-choice one for two reasons: 1) The open-ended task is more challenging as it requires models to choose from a large answer set rather than selecting from several given options. 2) Limiting the answer choices to only several options would significantly decrease the diversity of the dataset. Specifically, all answers in the QA pairs form an answer pool, which is treated as 191 classes (after discarding classes with fewer than 30 samples).
Why not Fine-Grained Action Recognition? It is essential to underscore that merely achieving fine-grained action recognition is insufficient for Sports-QA. The nature of the questions in this context demands not only precise detection of actions but also involves intricate temporal and causal modeling. To illustrate, questions that inquire about the effects of certain actions necessitate the accurate identification of temporally adjacent actions and a nuanced understanding of their causality. In summary, Sports-QA requires a more advanced level of comprehension and temporal reasoning beyond basic action recognition.